The area of a rectangle of length x is given by 15x-x2. Find the width of the rectangle in terms of x.
Thank you for your help in solving this problem.
Daniel H. Jack

Answer by Alan3354(69443)

The area of a rectangle of length x is given by 15x-x2. Find the width of the rectangle in terms of x.
----------
Area = L*W
(15x-x^2) = x*W
W = (15x-x^2)/x
W = 15 - x